OpenStudy (anonymous):

The perimeter of a rectangle is 56 inches. The length is 4 more than 3 times the width. What are the dimensions of the rectangle? Write the system of equations that would be used to solve this problem

OpenStudy (anonymous):

Perimeter of a rectangle: 2l+2w=56 Since the length is 4 more than 3 times the width, then l=3w+4 Thus, \[P=2(3w+4)+2w=56\] \[6w+8+2w=56\] \[8w+8=56\] Then:\[8w=48\] \[w=6\] Then plug it back into the length equation to find length: \[l=3(6)+4\] \[l=22\] Now width is 6 inches. If the length is 4 more than 3 times then width, then it should be 22, which is what I got.

OpenStudy (anonymous):

L=3w+4 P=2(w+L)=2(w+3w+4)=56 8w+8=56